26. A builder is adding a 20' × 28' room to an existing house. The
flooring choices are an 8" × 8" tile that costs $0.98 per tile, or a
12" × 12" tile that costs $1.98 per tile. Approximately how much
will the builder save by using the less expensive option?

Answer by checkley71(8403)

20*28=560 square feet.
the 8x8 inch tile is 64 square inches or 64/144=4/9 square yards. Therefore it will take
560/(4/9) tiles to tile this addition
(560*9)/4=1260 8x8 tiles are needed
1260*.98=$1,234.80 is the related cost for this type of tile.
---------------------------------------------------------------------
12*12 inch tiles = 1 square foot. Thus it will take
560/1=560 tiles @ a cost of 1.98. thus
560*1.98=$1,108.80 for this type of tile.
---------------------------------------------------------------------
1234.80-1108.80=$126.00 so the 12x12 inch tile is a better buy by $126.00.